# arXiv:2603.02196v1[cs.AI] 02 Mar 2026

### v1

@@ 4 Methods and Theory
Figure 4: Rejection sampling for policy interpolation at three $\beta$ values. As $\beta$ increases from near-zero to large values, the constrained policy $\pi^{\beta}$ interpolates from the safe policy $\pi_{0}$ to the optimized policy $\pi_{t}$. Blue histogram bars show accepted samples matching $\pi^{\beta}$, while teal/red bars show rejected samples from the proposal distribution. (a) Small $\beta$ nearly recovers the safe policy (b) Intermediate $\beta$ shows intermediate interpolation with frequent rejections. (c) Large $\beta$ approaches the optimized policy with minimal rejection.
@@ 4.1 Extending Conformal Risk Control to Non-Monotonic Losses
when the set is nonempty, and otherwise define $\widehat{\lambda}_{+}(\lbag L_{1:m}\rbag,\alpha):=\lambda_{\max}$, where a $\lambda_{\max}$ is a “safe hyperparameter” that we assume exists as does Angelopoulos et al. (2024): $\lambda_{\max}:=\sup\Lambda\in\Lambda$ and $L_{i}(\lambda_{\max})\leq\alpha$. Note that $\widehat{\lambda}_{+}(\lbag L_{1:m}\rbag,\alpha)$ can roughly be interpreted as the smallest (most aggressive) hyperparameter such that the empirical risk, $\sum_{i=1}^{m}\frac{L_{i}(\lambda)}{m}$, is controlled at level $\alpha$ for all hyperparameters at least as large as it. That is, the “for all $\lambda\geq\lambda_{0}$” qualification in Eq. (2) is the key algorithmic modification to standard CRC. We then define the oracle solution by Eq. (2) fit on the bag (or multiset) containing both the calibration and test loss functions, $\lbag L_{1:n+1}\rbag$:
@@
$\mathbb{E}\Big[\big|h(\lbag Z_{1:n+1\backslash i})-h(\lbag z_{1:n+1\backslash j})\big|\Big]\leq\epsilon.$
@@
The $\hat{\lambda}_{+}$ in Eq. (4) can be viewed as the smallest hyperparameter such that the “monotized empirical risk” is controlled at $\alpha$. The procedure is presented with this interpretation in Angelopoulos et al. (2024)’s App. A, but that paper noted that finite-sample guarantees for $\hat{\lambda}_{+}$ were at the time unknown.
@@
  $\pi_{t}^{(\beta)}(A_{i})\leftarrow\min(\rho_{i},\beta)\,/\,\psi_{\beta}$ for $i\in\mathcal{D}_{\text{cal}}$
@@ 4.2 Conformal Policy Control: Selecting a Risk-Controlling Policy
Assumption 1: An Initial Safe Policy Assume access to an initial safe policy, $\pi_{0}$, for which the true risk is controlled at the desired level $\alpha$; that is, assume $\mathbb{E}_{X\sim p}\ \mathbb{E}_{A\sim\pi_{0}(\cdot\mid X)}\ [\ell(X,A)]\leq\alpha$ holds for $\pi_{0}$.
@@
$\pi_{t}^{(\beta)}(x)\propto\min\big(\pi_{t}(x),\ \beta\cdot\pi_{0}(x)\big).$ (5)
Assume there is some small, positive $\beta_{\min}\ll 1$ such that $\beta_{\min}\leq\pi_{t}(x)/\pi_{0}(x)$ for all $x\in\mathcal{X}$. Then, note that as $\beta\to\beta_{\min}$, the constrained policy approaches the safe one: $\pi_{t}^{(\beta)}\to\pi_{0}$. Also note that as $\beta\to\infty$, the constrained policy approaches the optimized: $\pi_{t}^{(\beta)}\to\pi_{t}$ (Figure 4).
Robot Arm Kinematics
Airfoil
Healthcare Utilization (MEPS)
Calibrating $\boldsymbol{\beta}$: Like $\lambda$ in Section 4.1, $\beta$ is a control parameter, but it acts on the policy rather than the loss. We calibrate $\beta$ as we fit $\hat{\lambda}_{+}$, by searching from what we a priori expect is safest to most aggressive. Here for $\beta$, this means searching in increasing order in a grid (or set of intervals) $G\subset(0,\infty]$, to find the largest value where the conservatively-adjusted weighted empirical risk is still controlled:
@@
where the $\tilde{w}_{i}^{(\beta)}$ and $\tilde{w}_{\max}^{(\beta)}$ are carefully constructed (normalized) conformal importance weights. That is, for our theory, we define $\tilde{w}_{i}^{(\beta)}$ as in Prinster et al. (2024) via computing the likelihoods of all permutations of the calibration data $z_{0:t-1}$ and a hypothetical test sample $z_{t}\in\mathcal{Z}:=\mathcal{X}\times\mathcal{A}$, but a key subtlety here is that $z_{t}$ is not yet known, as CPC calibration is prior to sampling. For any $z_{t}$, define unnormalized weight
@@
$\tilde{w}_{i}^{(\beta)}:=w_{i}^{(\beta)}/\sum_{i=0}^{t}w_{i}^{(\beta)}$
and the conservative weight for an unknown test point as
Hereon, we refer to normalized calibration and (conservative) test point weights, that is, assuming $\sum_{i=0}^{t-1}\tilde{w}_{i}^{(\beta)}+\tilde{w}_{\max}^{(\beta)}=1$. In practice these weights are not computationally tractable, but we estimate the $\tilde{w}_{i}^{(\beta)}$ by approximating past policies as one mixture distribution, $\pi_{0:t-1}^{\text{mix}}\approx\pi_{0:t-1}$, which yields a fast linear runtime; additionally, we estimate $\tilde{w}_{\max}^{(\beta)}$ by an empirical maximum over samples from $\pi_{0}$ and $\pi_{t}$. If the weights are computed exactly, CPC achieves the following guarantee.
@@
Note that unlike Theorem 4.2, where the Lipschitz condition is on $l_{i}(\lambda)$, here the Lipschitz condition is on $w_{i}^{(\beta)}$ or $(l_{i}-\alpha)\cdot w_{i}^{(\beta)}$ as a function of $\beta$. Since $l_{i}$ does not depend on $\beta$, this regularity is determined entirely by the conformal weights, which depend on the likelihood ratio between the safe and optimized policies. The assumption thus concerns the agent’s own policies rather than the unknown constraint function.
Sampling from the Constrained Policy in Combinatorial Action Spaces In large action spaces, the normalization constant of $\pi_{t}^{(\hat{\beta})}$ is intractable. However, accept-reject (AR) sampling (Von Neumann and others, 1963) avoids normalization entirely. When $\hat{\beta}$ is small, proposals from $\pi_{0}$ are efficient, with the AR envelope constant set to $\hat{\beta}$. When $\hat{\beta}$ is large, proposals from $\pi_{t}$ are more efficient, since $\pi_{t}^{(\hat{\beta})}\approx\pi_{t}$. See Figure 4 and Appendix C.2 for details.
@@ 5.1 Factual Medical Question-Answering
Setup: We evaluate CRC methods on the MedLFQA dataset (Jeong et al., 2024), which aggregates medical question-answering benchmarks. We use the dataset of Cherian et al. (2024), which contains GPT-3.5-Turbo responses, atomic subclaims extracted via GPT-4o, and factuality annotations obtained by checking each subclaim against reference answers. Given a confidence score $p(C_{ij}\mid X_{i})$ for each subclaim, we filter claims below a threshold $\tau$ and measure the false discovery rate (FDR), the fraction of retained claims that are false, as our loss. We measure recall, the fraction of total true claims retained, as our utility metric. Unlike the binary loss used by Mohri and Hashimoto (2024) and Cherian et al. (2024), FDR is non-monotonic in $\tau$, making it a natural test case for gCRC. We compare gCRC against monotonized-losses CRC and Learn Then Test (LTT) (Angelopoulos et al., 2025). See Appendix D for details.
Results: Figure 5 shows FDR and recall across varying risk levels $\alpha$. gCRC controls FDR at or below the target level across all $\alpha$ values while achieving recall at least as high, or often significantly higher, than both the monotonized-losses CRC and LTT baselines. Relative to gCRC, the conservatism of LTT is due to its use of concentration inequalities for a high-probability guarantee, and that of monotized-losses CRC is a consequence of Jensen’s inequality (monotizing losses is far more conservative than monotizing the empirical risk or average loss value). These results demonstrate that gCRC provides valid risk control for non-monotonic losses where standard CRC may not be valid, and empirically with higher power than baselines.
@@ 6 Discussion
We set out to resolve a circularity. The safe policy’s deployment data can be reweighted to estimate the risk of a new policy, but the correct weights depend on which policy is deployed, and the deployed policy depends on the weighted risk estimates. Conformal policy control resolves this by parameterizing the deployed policy as a likelihood-ratio threshold, then calibrating that threshold on the safe policy’s data to find the most aggressive setting that still respects the risk threshold.
@@
Safety and exploration need not run counter to each other. In the right balance, they correct each other’s deficiencies, avoiding complacency on the one hand and dissipation on the other. However, it must be possible to determine an acceptable balance from what we actually know, not what we hope to be true. It can be enough, it turns out, to know what works, and to know what we want to do next.
@@ Acknowledgments
D.P, A.L., and S. Saria were partially funded by the Gordon and Betty Moore Foundation grant #12128. The authors would like to thank Hannah Lawrence and Neha Verma for helpful feedback.

### v2

@@ 4.1 Extending Conformal Risk Control to Non-Monotonic Losses
when the set is nonempty, and otherwise define $\widehat{\lambda}_{+}(\lbag L_{1:m}\rbag,\alpha):=\lambda_{\max}$, where $\lambda_{\max}$ is a “safe hyperparameter” assumed to exist as in Angelopoulos et al. (2024): $\lambda_{\max}:=\sup\Lambda\in\Lambda$ and $L_{i}(\lambda_{\max})\leq\alpha$. In words, $\widehat{\lambda}_{+}$ is the smallest (most aggressive) hyperparameter such that the empirical risk is controlled at level $\alpha$ for all hyperparameters at least as large as it; this “for all $\lambda\geq\lambda_{0}$” qualification is the key algorithmic modification to standard CRC. We then define the oracle solution by Eq. (2) fit on the bag (or multiset) containing both the calibration and test loss functions, $\lbag L_{1:n+1}\rbag$:
@@
$\mathbb{E}\Big[\big|h(\lbag Z_{1:n+1}\rbag\backslash\{z_{i}\})-h(\lbag Z_{1:n+1}\rbag\backslash\{Z_{j}\})\big|\Big]\leq\epsilon.$
@@
The $\hat{\lambda}_{+}$ in Eq. (4) can be viewed as the smallest hyperparameter such that the “monotonized empirical risk” is controlled at $\alpha$. The procedure is presented with this interpretation in Angelopoulos et al. (2024)’s App. A, but that paper noted that finite-sample guarantees for $\hat{\lambda}_{+}$ were at the time unknown.
@@
  $\pi_{t}^{(\beta)}(A_{i})\leftarrow\min(\pi_{t}(A_{i}),\beta\cdot\pi_{0}(A_{i}))\,/\,\psi_{\beta}$ for $i\in\mathcal{D}_{\text{cal}}$
@@ 4.2 Conformal Policy Control: Selecting a Risk-Controlling Policy
Assumption 1: An Initial Safe Policy Assume access to an initial safe policy, $\pi_{0}$, for which the true risk is controlled at the desired level $\alpha$; that is, assume $\mathbb{E}_{X\sim p}\ \mathbb{E}_{A\sim\pi_{0}(\cdot\mid X)}\ [\ell(X,A)]\leq\alpha$ holds for $\pi_{0}$. Note that this assumption is weaker than it may appear: any existing policy (a previous deployment, a domain default, or even a uniform baseline) qualifies as $\pi_{0}$ so long as its risk is at most $\alpha$. When little is known, $\alpha$ can be set loosely and any reasonable baseline suffices; the strength of the resulting guarantee scales with how much is actually known about what is safe.
@@
$\pi_{t}^{(\beta)}(a\mid x)\propto\min\big(\pi_{t}(a\mid x),\ \beta\cdot\pi_{0}(a\mid x)\big).$ (5)
Assume there is some small, positive $\beta_{\min}\ll 1$ such that $\beta_{\min}\leq\pi_{t}(a\mid x)/\pi_{0}(a\mid x)$ for all $x\in\mathcal{X}$, $a\in\mathcal{A}$. As $\beta\to\beta_{\min}$ the constrained policy approaches the safe one, $\pi_{t}^{(\beta)}\to\pi_{0}$, and as $\beta\to\infty$ it approaches the optimized one, $\pi_{t}^{(\beta)}\to\pi_{t}$ (Figure 4). Crucially, because $\beta$ governs the likelihood ratio and not the loss, any smoothness needed for a guarantee will be smoothness the agent can compute and verify. Theorem 4.5 makes this precise.
Calibrating $\boldsymbol{\beta}$: We calibrate $\beta$ analogously to $\hat{\lambda}_{+}$ (Section 4.1), searching from what we a priori expect is safest to most aggressive. For $\beta$, this means searching in increasing order in a grid (or set of intervals) $G\subset(0,\infty]$, to find the largest value where the conservatively-adjusted weighted empirical risk is still controlled (Algorithm 1):
@@
where the $\tilde{w}_{i}^{(\beta)}$ and $\tilde{w}_{\max}^{(\beta)}$ are (normalized) conformal importance weights. These weights correct for the distribution shift between the policies that generated the calibration data and the candidate policy $\pi_{t}^{(\beta)}$: since the agent’s past actions were drawn from earlier policies, each observation must be reweighted to obtain a valid risk estimate under $\pi_{t}^{(\beta)}$. This is the importance weighting described in Section 2, now made precise for the multi-round setting.
Formally, we define $\tilde{w}_{i}^{(\beta)}$ as in Prinster et al. (2024) via the likelihoods of all permutations of the calibration data $z_{0:t-1}$ and a hypothetical test sample $z_{t}\in\mathcal{Z}:=\mathcal{X}\times\mathcal{A}$. A key subtlety is that $z_{t}$ is not yet known, as CPC calibration occurs prior to sampling. For any $z_{t}$, define the unnormalized weight
@@
$\tilde{w}_{i}^{(\beta)}:=w_{i}^{(\beta)}/\sum_{i=0}^{t}w_{i}^{(\beta)},\hskip 11.38109pt\tilde{w}_{\max}^{(\beta)}:=\sup_{z_{t}\in\mathcal{Z}}(\tilde{w}_{t}^{(\beta)}),$
where $\tilde{w}_{\max}^{(\beta)}$ is a conservative estimate of the test point weight. Hereon, we refer to normalized calibration and (conservative) test point weights, that is, assuming $\sum_{i=0}^{t-1}\tilde{w}_{i}^{(\beta)}+\tilde{w}_{\max}^{(\beta)}=1$. In practice these weights are not computationally tractable, but we estimate the $\tilde{w}_{i}^{(\beta)}$ by approximating past policies as one mixture distribution, $\pi_{0:t-1}^{\text{mix}}\approx\pi_{0:t-1}$, which yields a fast linear runtime; additionally, we estimate $\tilde{w}_{\max}^{(\beta)}$ by an empirical maximum over samples from $\pi_{0}$ and $\pi_{t}$. If the weights are computed exactly, CPC achieves the following guarantee.
@@
A key advantage of CPC over gCRC is where the Lipschitz condition falls. In Theorem 4.2, smoothness is required of $l_{i}(\lambda)$, the loss as a function of the control parameter, which is a property of the unknown environment. In Theorem 4.5, smoothness is required of $w_{i}^{(\beta)}$ or $(l_{i}-\alpha)\cdot w_{i}^{(\beta)}$ as a function of $\beta$. Since $l_{i}$ does not depend on $\beta$, this regularity is determined entirely by the conformal weights, which depend on the likelihood ratio between the safe and optimized policies. The assumption thus concerns the agent’s own policies rather than the unknown constraint function, and is in principle verifiable or enforceable by the practitioner.
Sampling from the Constrained Policy in Combinatorial Action Spaces In large action spaces, the normalization constant of $\pi_{t}^{(\hat{\beta})}$ is intractable. However, accept-reject (AR) sampling (Von Neumann and others, 1963) avoids normalization entirely. When $\hat{\beta}$ is small, proposals from $\pi_{0}$ are efficient, with the AR envelope constant set to $\hat{\beta}$. When $\hat{\beta}$ is large, proposals from $\pi_{t}$ are more efficient, since $\pi_{t}^{(\hat{\beta})}\approx\pi_{t}$. See Figure 4 for an illustration and Appendix C.2 for details.
Robot Arm Kinematics
Airfoil
Healthcare Utilization (MEPS)
@@ 5.1 Factual Medical Question-Answering
Setup: We evaluate CRC methods on the MedLFQA dataset (Jeong et al., 2024), which aggregates medical question-answering benchmarks. We use the dataset of Cherian et al. (2024), which contains GPT-3.5-Turbo responses, atomic subclaims extracted via GPT-4o, and factuality annotations obtained by checking each subclaim against reference answers. Given a confidence score $p(C_{ij}\mid X_{i})$ for each subclaim, we filter claims below a threshold $\tau$ and measure the false discovery rate (FDR), the fraction of retained claims that are false, as our loss. We measure recall, the fraction of total true claims retained, as our utility metric. Unlike the binary loss used by Mohri and Hashimoto (2024) and Cherian et al. (2024), FDR is non-monotonic in $\tau$, making it a natural test case for gCRC. We compare gCRC against CRC applied to monotonized losses and Learn Then Test (LTT) (Angelopoulos et al., 2025). See Appendix D for details.
Results: Figure 5 shows FDR and recall across varying risk levels $\alpha$. gCRC controls FDR at or below the target level across all $\alpha$ values while achieving recall at least as high, or often significantly higher, than both the monotonized-losses CRC and LTT baselines. Relative to gCRC, LTT is conservative because it uses concentration inequalities to obtain a high-probability guarantee. Monotonized-loss CRC is conservative as a consequence of Jensen’s inequality: monotonizing each loss is far more conservative than monotonizing the empirical risk. These results demonstrate that gCRC provides valid risk control for non-monotonic losses where standard CRC may not be valid, and empirically with higher power than baselines.
@@ 6 Discussion
We set out to let an agent choose how far to deviate from a safe policy without the act of choosing biasing the estimate that justified the choice. The resolution exploits the fact that the only shift between the safe policy’s data and any candidate’s deployment is the agent’s own: it does not know how the world will respond, but it knows exactly how each candidate would distribute its actions. Conformal policy control parameterizes that choice as a likelihood-ratio threshold and calibrates it directly on the safe policy’s data, so that selection is the calibrated act rather than a step that follows it. Because the threshold bounds how far the candidate can deviate from the safe policy, the resulting guarantee applies to the chosen policy itself, not merely each candidate in isolation.
More broadly, CPC instantiates a prescriptive use of conformal methods: rather than characterizing uncertainty under a fixed distribution, it searches over distributions to find one that satisfies a finite-sample guarantee. This inversion applies wherever one selects a data-generating process subject to a risk constraint, from experimental design, to active data collection, to generative model deployment. We hope that the ideas in this paper encourage further development of conformal methods as tools for decision-making, beyond their traditional role in uncertainty quantification.
@@
A recurring theme in this work is that assumptions should reflect what the practitioner knows rather than what is theoretically convenient. CPC requires a safe policy, but any baseline whose risk the practitioner is willing to accept will do. It requires smoothness, but of the agent’s own policy likelihood ratios, not of the unknown constraint loss function. It requires a risk tolerance, but as a direct declaration rather than an indirect penalty weight. Safety and exploration, in this view, need not run counter to each other. In the right balance they correct each other’s deficiencies, avoiding complacency on the one hand and dissipation on the other. The balance must be set from what is actually known, not what is hoped. It can be enough, it turns out, to know what works, and to know what we want to do next.
@@ Acknowledgments
D.P., A.L., and S. Saria were partially funded by the Gordon and Betty Moore Foundation grant #12128. The authors would like to thank Hannah Lawrence and Neha Verma for helpful feedback.